1. Todays stock price reflects investor expectations about the earning power of the firms current and future assets. Take Google, for example. All its earnings are plowed back into new invest ments and the stock sells at price of $344. Suppose that the earnings from Googles existing business are expected to stay constant in real terms. Investors are valuing Googles future investment opportunities at $164. Find Googles current earnings per share. (Investors also expect an estimated 7.4% real cost of equity) (a) $180 (b) $34 (c) $13 2. Machines A and B are mutually exclusive and are expected to produce the following real cash flows. Calculate the equivalent annual cash flow from each machine and decide which machine you should buy. Assume the cost of capital is 4%. Cash Flows ($ thousands) C 2 +121 +121 Machine +110 -100 120 +110 +133 (a) A: 62.37/B: 77.79 / Choose B (b) A: 49.54 /B: 59.41/ Choose A (c) A: 33.54 / B: 37.46/ Choose B (d) A: 16.98/ B: 24.87/ Choose A
Answer #1

1. Current earnings per share = ( Price of stock * Cost of equity ) - ( Opportunity of future investments * Cost of equity)

= (344*7.4% ) - (164*7.4%)

= 13 per share

Current earnings per share = 13 per share

2. The formula for calculating the equivalent annual cash flows is

              C = [r ( NPV)] / [1 - (1+ r)^(-n)]

NPV for machine A = -100 + 110 /( 1+ 0.04)1 + 121/ ( 1+ 0.04)2 = 117.64

NPV for machine B = -120 + 110 /( 1+ 0.04)1 + 121/ ( 1+ 0.04)2 + 133 / ( 1+0.04)3 = 215.88

equivalent annual cash flows

For Machine-A:

                                  C = [0.04* $117.64] / [1-(1+0.04)-2]

                                     = $4.70 / 0.0754

                                     = $62.3

For Machine-B:

                                C = [ 0.04 * $215.88] / [1 - (1 + 0.04)-3]

                                   = $8.6352 / 0.111

                                   = $77.79

By Companring the two projects, the Machine-B with 3yrs is yielding highest NPV and returning high amount of annuity cash flows relative to the other Machine-A. Hence, project-B should be selected.

So ans is a) A: 62.3 / B: 77.79 / choose B